OpenStudy (anonymous):

HELP!!!! The bus fare in a city is $1.50. People who use the bus have the option of purchasing a monthly coupon book for $24.00. With the coupon book, the fare is reduced to $0.50. Determine the number of times in a month the bus must be used so that the total monthly cost without the coupon book is the same as the total monthly cost with the coupon book.

OpenStudy (anonymous):

Let x = number of bus fares. Cost of fares without coupon book = 1.50x Cost of fares WITH coupon book = .50x + 24 1.50x = .50x+24 To solve the equation, subtract .50x from each side of the equation: 1.50x -.50x = .50x-.50x + 24 1.00x = 24 x= 24 fares

OpenStudy (anonymous):

x = number of times you ride on the bus y = total cost without book: y = 1.5x with book: y = 24 + 0.5x Using systems of equations (equal values method), set the equations equal to each other 1.5x = 24+0.5x Solve with algebra x = 24 times on the bus